OpenStudy (anonymous):

want you to build me a pyramid out of plywood. the side of the pyramid will be EQUILATERAL TRAINGLES with sides measuring 8ft. what is the LATERAL AREA of the pyramid? round to the nearest ft. PLEASE HELP ME, im homeschooled

OpenStudy (cwrw238):

the area of each triangular side will be 0.5 * base*height 

OpenStudy (cwrw238):

the height can be found using the pythagoras theorem: h^2 = 8^2 - 4^2 # 56 h = sqrt56 = 7.48 ft area = 0.5 * 8 * 7.48 = 29.933 sq ft

OpenStudy (cwrw238):

you dont say what shape the base of the triangle is. if, for example its square you multiply the above area by 4, or if its a triangular base multiply by 3